# Learning-based Distributed Model Predictive Control using Multi-Agent Bayesian Optimization

### v1

@@ III-C Parameterized DMPC
In this paper, we propose to use a parameterized DMPC scheme as approximators for the local policy $\boldsymbol{\mathrm{\pi}}_{i},i=1,\ldots,m$. More precisely, we aim to learn a local MPC parameterized by $\boldsymbol{\mathrm{\theta}}_{i}$ such that the parameterized policy $\boldsymbol{\mathrm{\pi}}_{\boldsymbol{\mathrm{\theta}}_{i}}$ delivered from each local MPC can provide an accurate approximation of the optimal policy $\boldsymbol{\mathrm{\pi}}_{i}^{\star}$ at $\boldsymbol{\mathrm{\zeta}}^{\star}:=\left\{\boldsymbol{\mathrm{\theta}}_{1}^{\star},\ldots,\boldsymbol{\mathrm{\theta}}_{m}^{\star}\right\}$ even if the MPC model cannot capture the real system perfectly. By modification of the DMPC scheme, the following then holds:
@@
To tackle the local optimization problem (19) at every time instant $k$, we use a parameterized DMPC based on dual-decomposition. We then propose to parameterize each local MPC scheme so that the local value function $\hat{V}_{i}^{N,\boldsymbol{\mathrm{\pi}}}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k}\right)$ can be replaced by $V_{i}^{\boldsymbol{\mathrm{\theta}}_{i}}$ delivered from the local MPC. To capture the local optimal policies $\boldsymbol{\mathrm{\pi}}_{i}^{\star}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k}\right)$, one needs to learn the corresponding local MPC schemes in a coordinated manner such that $\boldsymbol{\mathrm{\zeta}}:=\left\{\boldsymbol{\mathrm{\theta}}_{1},\ldots,\boldsymbol{\mathrm{\theta}}_{m}\right\}\rightarrow\boldsymbol{\mathrm{\zeta}}^{\star}$ results in
@@
Taking into account the previous observation in Remark 3, the value function captured from a parametrized local MPC scheme must be a function of the total state $\boldsymbol{\mathrm{\tilde{s}}}^{k}$. Although the value function associated to (5) is implicitly a function of $\boldsymbol{\mathrm{\tilde{s}}}^{k}$ through the multipliers $\boldsymbol{\mathrm{\mu}}_{i}$ and the slack variables $\bar{\boldsymbol{\mathrm{w}}}_{i}^{\ell}$, it is only a function of the local state $\boldsymbol{\mathrm{s}}_{i}^{k}$ explicitly. Hence, the same modification as Theorem 1 will not work for the DMPC scheme (5). To address this issue, without loss of generality, we propose to force the initial slack variables to be $\bar{\boldsymbol{\mathrm{w}}}_{i}^{k}=W_{ij}\left({\boldsymbol{\mathrm{s}}}_{j}^{k}\right)$. We also introduce an additional constraint to smooth the slack variables. This constraint also acts as a model for propagating the effect of $W_{ij}\left(\ {\boldsymbol{\mathrm{s}}}_{j}^{k}\right)$ along the prediction horizon. The proposed DMPC scheme is then formulated as
@@
$\qquad\boldsymbol{\mathrm{h}}_{i}\left(\hat{\boldsymbol{\mathrm{x}}}_{i}^{\ell},\hat{\boldsymbol{\mathrm{u}}}_{i}^{\ell}\right)\leq 0,\quad\boldsymbol{\mathrm{h}}_{i}\left(\hat{\boldsymbol{\mathrm{x}}}_{i}^{k+N}\right)\leq 0,$ (22e)
We next show that the value function associated to the proposed DMPC scheme above can capture the true value function $V_{i}^{\boldsymbol{\mathrm{\pi}}}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k}\right)$ even if the underlying local models $\boldsymbol{\mathrm{f}}_{i}$ cannot capture the real subsystems (agents).
@@
Considering Remark 1 and under Assumption 2, one can find the modified cost functions $\hat{T}_{i}^{\boldsymbol{\mathrm{\mu}}^{(I)}}$ and $\hat{L}_{i}^{\boldsymbol{\mathrm{\mu}}^{(I)}}$ such that the value function associated to (22) can capture $V_{i}^{\boldsymbol{\mathrm{\pi}}}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k}\right)$ at $\boldsymbol{\mathrm{\mu}}^{\star}$ even if the trajectories are imperfectly evolved by $\boldsymbol{\mathrm{\hat{f}}}_{i}$.
@@
Assuming that the trajectories $\hat{\boldsymbol{\mathrm{s}}}^{1},\ldots,\hat{\boldsymbol{\mathrm{s}}}^{N}$ in (14) are approximately deterministic, one can adopt the same structure without expectation to define the value function associated to (22) with a possibly wrong deterministic model. The modified value function is then defined as
@@
$\qquad\qquad\qquad\qquad\qquad+\sum_{\ell=k}^{k+N-1}\gamma^{\ell-k}\hat{L}_{i}^{\boldsymbol{\mathrm{\mu}}^{\star}}\left(\hat{\boldsymbol{\mathrm{x}}}_{i}^{\ell},\hat{\boldsymbol{\mathrm{u}}}_{i}^{\ell},\bar{\boldsymbol{\mathrm{w}}}_{i}^{\ell}\right).$
where $\hat{T}_{i}^{\boldsymbol{\mathrm{\mu}}^{\star}},\hat{L}_{i}^{\boldsymbol{\mathrm{\mu}}^{\star}}$ are the modified version of the costs in (22) with the last argument dropped. We then select the modified terminal and stage costs same as Theorem 1 as
@@
Under Remark 1 and applying the initial conditions $\hat{\boldsymbol{\mathrm{x}}}_{i}^{k}=\boldsymbol{\mathrm{s}}_{i}^{k}$ and $\bar{\boldsymbol{\mathrm{w}}}_{i}^{k}=W_{ij}\left({\boldsymbol{\mathrm{s}}}_{j}^{k}\right)$, we then have that $\hat{V}^{\text{DMPC}}=V_{i}^{\boldsymbol{\mathrm{\pi}}}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k}\right)$. ∎
By Theorem 1 and its Corollary 1, we prove that the DMPC cost function can be modified such that the local value functions associated to the true multi-agent MDP are captured by the modified local MPC schemes. More specifically, the central theorem aims to show that there exists such a modification and to understand its structure. However, the proposed modification structure is not tractable in terms of implementation. To tackle this problem, we propose to parameterize the corresponding cost terms and learn them in a coordinated manner for the best closed-loop performance. We then propose to parameterize a discounted version of the DMPC scheme (22) at $\hat{\boldsymbol{\mathrm{\mu}}}^{\star}$ as
@@
IV Fusion of MABO and DMPC
We assume that the dynamics $\boldsymbol{\mathrm{x}}_{i}^{k+1}=\boldsymbol{\mathrm{f}}_{i}(\boldsymbol{\mathrm{x}}_{i}^{k},\,\boldsymbol{\mathrm{u}}_{i}^{k}),i=1,\ldots,m$ cannot capture the real multi-agent system exactly, such that there is a model-mismatch between the real system and the local MPC models dedicated to each agent. To account for this mismatch, we leverage the results detailed in Section III to modify the cost functions of the local MPC schemes (and possibly their models and constraints) in a coordinated manner leading to the best closed-loop performance of each agent, despite the model-mismatch. To this end, we employ an ADMM-based multi-agent Bayesian optimization proposed in [22] and infer the optimal parameterization from closed-loop data.
IV-A Background on Bayesian Optimization
Bayesian Optimization (BO) is an effective approach for optimizing complex, expensive, and noisy black-box functions. In control applications, it is particularly useful for learning parameters in controller such as MPC. However, there are no existing works exploiting the BO methods for learning networked model-based control systems such as distributed MPC where the local models cannot represent the true multi-agent system.
The BO methods use Gaussian Process (GP) surrogate models to learn an approximation of the not explicitly known function with respect to some parameters, i.e., the closed-loop performance of a parameterized MPC $J\left(\boldsymbol{\mathrm{\theta}}\right):=J\left(\boldsymbol{\mathrm{\pi}}^{\boldsymbol{\mathrm{\theta}}}\right)$ with respect to the policy parameters $\boldsymbol{\mathrm{\theta}}$. Considering $J:\mathbb{R}^{n_{\boldsymbol{\mathrm{\theta}}}}\rightarrow\mathbb{R}$, we then define the corresponding GP model as
@@ IV-A Background on Bayesian Optimization
where $\mu(\boldsymbol{\mathrm{\theta}})=\mathbb{E}[J(\boldsymbol{\mathrm{\theta}})]$ is the mean function, typically assumed to be zero, and the kernel $k(\boldsymbol{\mathrm{\theta}},\boldsymbol{\mathrm{\theta}}^{\prime})$ is the covariance function, determining the relationship between points in the input space. More precisely, the covariance function describes how the values of the function relate to each other. A commonly used kernel is the squared exponential kernel:
@@
where where $\sigma_{J}^{2}$ is the signal variance, and $l$ is the length scale. The choice of kernel significantly impacts the behavior of the GP model.
Given observed data $\mathcal{D}=\{(\boldsymbol{\mathrm{\theta}}^{k},y^{k})\}_{k=1}^{n}$, where $y^{k}=J(\boldsymbol{\mathrm{\theta}}^{k})+\epsilon$ and $\epsilon\sim\mathcal{N}(0,\sigma_{n}^{2})$:
@@
where $\boldsymbol{\mathrm{\Theta}}=\left\{\boldsymbol{\mathrm{\theta}}^{k}\right\}_{k=1}^{n}$ and $K$ denotes the covariance matrix ($n\times n$ kernel matrix with elements $[K]_{(i,j)}=k(\boldsymbol{\mathrm{\theta}}_{i},\boldsymbol{\mathrm{\theta}}_{j})$) computed over the training inputs. The posterior distribution of $f(\boldsymbol{\mathrm{\theta}}_{\star})$ at a new input $\boldsymbol{\mathrm{\theta}}_{\star}$ is then given by:
@@
$\begin{bmatrix}\mathbf{y}\\ J_{\star}\end{bmatrix}\sim\mathcal{N}\left(\begin{bmatrix}\mu(\boldsymbol{\mathrm{\Theta}})\\ \mu(\boldsymbol{\mathrm{\theta}}_{\star})\end{bmatrix},\begin{bmatrix}K+\sigma_{n}^{2}I&k(\boldsymbol{\mathrm{\theta}}_{\star},\boldsymbol{\mathrm{\Theta}})\\ k(\boldsymbol{\mathrm{\Theta}},\boldsymbol{\mathrm{\theta}}_{\star})&k(\boldsymbol{\mathrm{\theta}}_{\star},\boldsymbol{\mathrm{\theta}}_{\star})\end{bmatrix}\right).$
@@
This allows us to derive the predictive mean and variance for $J(\boldsymbol{\mathrm{\theta}}_{\star})$:
$\mu(\boldsymbol{\mathrm{\theta}}_{\star})=k(\boldsymbol{\mathrm{\theta}}_{\star},\boldsymbol{\mathrm{\Theta}})(K+\sigma_{n}^{2}I)^{-1}\mathbf{y},$
$\sigma^{2}(\boldsymbol{\mathrm{\theta}}_{\star})=k(\boldsymbol{\mathrm{\theta}}_{\star},\boldsymbol{\mathrm{\theta}}_{\star})-k(\boldsymbol{\mathrm{\theta}}_{\star},\boldsymbol{\mathrm{\Theta}})(K+\sigma_{n}^{2}I)^{-1}k(\boldsymbol{\mathrm{\Theta}},\boldsymbol{\mathrm{\theta}}_{\star}).$
The Bayesian Optimization problem is formulated as
$\boldsymbol{\mathrm{\theta}}^{\star}=\arg\min_{\boldsymbol{\mathrm{\theta}}}J(\boldsymbol{\mathrm{\theta}}),$ (30)
The acquisition function $\alpha(\boldsymbol{\mathrm{\theta}})$ is crucial for guiding the optimization problem above. In this paper, we use the Expected Improvement (EI) as an acquisition function:
$\alpha_{\text{EI}}(\boldsymbol{\mathrm{\theta}})=\mathbb{E}[\max(J^{\star}-J(\boldsymbol{\mathrm{\theta}}),0)],$ (31)
@@
where $J^{\star}$ is the best observed function value. This acquisition function can be computed using the properties of the Gaussian distribution:
$Z=\frac{J^{*}-\mu(\boldsymbol{\mathrm{\theta}})}{\sigma(\boldsymbol{\mathrm{\theta}})},$
@@
$\alpha_{\text{EI}}(\boldsymbol{\mathrm{\theta}})=(J^{*}-\mu(\boldsymbol{\mathrm{\theta}}))\Phi(Z)+\sigma(\boldsymbol{\mathrm{\theta}})\phi(Z),$
where $\Phi(Z)$ and $\phi(Z)$ denote the standard normal Cumulative Density Function (CDF) and Probability Density Function (PDF) of the standard normal distribution, respectively.

### v2

@@ III-C Parameterized DMPC for Local Policy Approximation
In this section, we propose to use a parameterized DMPC scheme as approximators for the local policy $\boldsymbol{\mathrm{\pi}}_{i},i=1,\ldots,m$. More precisely, we aim to learn a local MPC parameterized by $\boldsymbol{\mathrm{\theta}}_{i}$ such that the parameterized policy $\boldsymbol{\mathrm{\pi}}_{\boldsymbol{\mathrm{\theta}}_{i}}$ delivered from each local MPC can provide an accurate approximation of the optimal policy $\boldsymbol{\mathrm{\pi}}_{i}^{\star}$ at $\boldsymbol{\mathrm{\zeta}}^{\star}:=\left\{\boldsymbol{\mathrm{\theta}}_{1}^{\star},\ldots,\boldsymbol{\mathrm{\theta}}_{m}^{\star}\right\}$ even if the MPC model cannot capture the real system perfectly. By modifying the DMPC scheme, we have
@@
To address the local optimization problem (25) at every time instant $k$, we use a parameterized DMPC based on dual decomposition. We then propose to parameterize each local MPC scheme so that the local value function $\hat{V}_{i}^{N,\boldsymbol{\mathrm{\pi}}}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k}\right)$ can be replaced by $V_{i}^{\boldsymbol{\mathrm{\theta}}_{i}}$ delivered from the local MPC. To capture the local optimal policies $\boldsymbol{\mathrm{\pi}}_{i}^{\star}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k}\right)$, one needs to learn the corresponding local MPC schemes in a coordinated manner such that $\boldsymbol{\mathrm{\zeta}}:=\left\{\boldsymbol{\mathrm{\theta}}_{1},\ldots,\boldsymbol{\mathrm{\theta}}_{m}\right\}\rightarrow\boldsymbol{\mathrm{\zeta}}^{\star}$ results in
@@
Taking into account the previous observation in Remark 3, the value function captured from a parametrized local MPC scheme must be a function of the complete state $\boldsymbol{\mathrm{\tilde{s}}}^{k}$. Although the value function associated with (5) is implicitly a function of $\boldsymbol{\mathrm{\tilde{s}}}^{k}$ through the multipliers $\boldsymbol{\mathrm{\mu}}_{i}$ and the slack variables $\bar{\boldsymbol{\mathrm{w}}}_{i}^{\ell}$, it is only a function of the local state $\boldsymbol{\mathrm{s}}_{i}^{k}$ explicitly. Hence, the same modification as in Theorem 1 will not work for the DMPC scheme (5). To address this issue, without loss of generality, we propose to force the initial slack variables to be $\bar{\boldsymbol{\mathrm{w}}}_{i}^{k}=W_{ij}\left({\boldsymbol{\mathrm{s}}}_{j}^{k}\right)$. We also introduce an additional constraint to smooth out the slack variables. This constraint also acts as a model for propagating the effect of $W_{ij}\left({\boldsymbol{\mathrm{s}}}_{j}^{k}\right)$ along the prediction horizon. The proposed DMPC scheme is then formulated as
@@
$\qquad\boldsymbol{\mathrm{h}}_{i}\left(\hat{\boldsymbol{\mathrm{x}}}_{i}^{\ell},\hat{\boldsymbol{\mathrm{u}}}_{i}^{\ell}\right)\leq 0,\quad\boldsymbol{\mathrm{h}}_{i}\left(\hat{\boldsymbol{\mathrm{x}}}_{i}^{k+N}\right)\leq 0.$ (28e)
We next show that the value function associated to the proposed DMPC scheme above can capture the true value function $V_{i}^{\boldsymbol{\mathrm{\pi}}}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k}\right)$ even if the underlying local models $\boldsymbol{\mathrm{f}}_{i}$ do not capture the real subsystems (agents).
@@
Considering Remark 1 and under Assumption 2, one can find the modified cost functions $\hat{T}_{i}^{\boldsymbol{\mathrm{\mu}}^{(I)}}$ and $\hat{L}_{i}^{\boldsymbol{\mathrm{\mu}}^{(I)}}$ such that the value function associated with (28) can capture $V_{i}^{\boldsymbol{\mathrm{\pi}}}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k}\right)$ at $\boldsymbol{\mathrm{\mu}}^{\star}$ even if the trajectories are imperfectly evolved by $\boldsymbol{\mathrm{\hat{f}}}_{i}$.
@@
Assuming that the trajectories $\hat{\boldsymbol{\mathrm{s}}}^{1},\ldots,\hat{\boldsymbol{\mathrm{s}}}^{N}$ in (15) are approximately deterministic, one can adopt the same structure without expectation to define the value function associated with (28) using a possibly wrong deterministic model. The modified value function is then defined as
@@
$\qquad\qquad\qquad\qquad\qquad+\sum_{\ell=k}^{k+N-1}\gamma^{\ell-k}\hat{L}_{i}^{\boldsymbol{\mathrm{\mu}}^{\star}}\left(\hat{\boldsymbol{\mathrm{x}}}_{i}^{\ell},\hat{\boldsymbol{\mathrm{u}}}_{i}^{\ell},\bar{\boldsymbol{\mathrm{w}}}_{i}^{\ell}\right),$
where $\hat{T}_{i}^{\boldsymbol{\mathrm{\mu}}^{\star}},\hat{L}_{i}^{\boldsymbol{\mathrm{\mu}}^{\star}}$ are the modified versions of the costs in (28) with the last argument dropped. We then select the modified terminal and stage costs to be the same as those in Theorem 1 as
@@
Under Remark 1 and applying the initial conditions $\hat{\boldsymbol{\mathrm{x}}}_{i}^{k}=\boldsymbol{\mathrm{s}}_{i}^{k}$ and $\bar{\boldsymbol{\mathrm{w}}}_{i}^{k}=W_{ij}\left({\boldsymbol{\mathrm{s}}}_{j}^{k}\right)$, we then obtain $\hat{V}^{\text{DMPC}}=V_{i}^{\boldsymbol{\mathrm{\pi}}}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k}\right)$. ∎
By Theorem 1 and Corollary 1, we prove that the DMPC cost function can be modified such that the local value functions associated with the true multi-agent MDP are captured by the modified local MPC schemes. More specifically, the central theorem aims at showing that there exists such a modification and at understanding its structure. However, the proposed modification structure is not tractable for implementation purposes. To address this, we propose to parameterize the corresponding cost terms and learn them in a coordinated manner for the best closed-loop performance. We then propose to parameterize a discounted version of the DMPC scheme (28) at $\hat{\boldsymbol{\mathrm{\mu}}}^{\star}$ as
@@
IV Fusion of Multi-agent Bayesian Optimization (MABO) and DMPC
We assume that the state-space model $\boldsymbol{\mathrm{x}}_{i}^{k+1}=\boldsymbol{\mathrm{f}}_{i}(\boldsymbol{\mathrm{x}}_{i}^{k},\,\boldsymbol{\mathrm{u}}_{i}^{k}),~{}i=1,\ldots,m$, cannot capture the true subsystems exactly, such that there is a model mismatch between the real system and the local MPC model each agent employs. To account for this mismatch, we leverage the results detailed in Section III to modify the cost function of the local MPC schemes (and possibly their models and constraints) in a coordinated manner leading to the optimal closed-loop performance of each agent despite the model mismatch. To this end, we employ an Alternating Direction Method of Multipliers (ADMM)-based multi-agent Bayesian optimization approach, initially proposed in [23], and infer the optimal parameterization from closed-loop data.
IV-A Review of Bayesian Optimization
Bayesian Optimization (BO) is an efficient approach for optimizing complex, expensive-to-obtain, and noisy black-box functions as
$\boldsymbol{\mathrm{\theta}}^{\star}=\arg\min_{\boldsymbol{\mathrm{\theta}}\in\mathcal{P}}J(\boldsymbol{\mathrm{\theta}}),$ (34)
where $\boldsymbol{\mathrm{\theta}}^{\star}$ is the global minimizer of an unknown objective function $J(\boldsymbol{\mathrm{\theta}})$ on the parameter space $\mathcal{P}$. In control applications, BO is particularly useful for learning control hyperparameters, e.g., in MPC. However, there are no existing works employing BO for learning networked model-based control schemes, such as distributed MPC, where the local models cannot represent the true multi-agent system. The BO methods use Gaussian Process (GP) surrogate models to learn an approximation of not-explicitly-known functions with respect to some parameters, e.g., the closed-loop performance of a parameterized MPC $J\left(\boldsymbol{\mathrm{\theta}}\right):=J\left(\boldsymbol{\mathrm{\pi}}^{\boldsymbol{\mathrm{\theta}}}\right)$ with respect to the policy parameters $\boldsymbol{\mathrm{\theta}}$. Considering $J:\mathbb{R}^{n_{\boldsymbol{\mathrm{\theta}}}}\rightarrow\mathbb{R}$, we then define the corresponding GP model as
@@ IV-A Review of Bayesian Optimization
where $\mu(\boldsymbol{\mathrm{\theta}})=\mathbb{E}[J(\boldsymbol{\mathrm{\theta}})]$ is the mean function, typically assumed to be zero, and the kernel $k(\boldsymbol{\mathrm{\theta}},\boldsymbol{\mathrm{\theta}}^{\prime})$ is the covariance function that relates points in the input space. A commonly used kernel is the squared exponential function as
@@
where $\sigma_{J}^{2}$ is the signal variance, and $l$ is the length scale. The choice of kernel significantly impacts the behavior of the GP model. Given observed data $\mathcal{D}^{k}=\{(\boldsymbol{\mathrm{\theta}}^{i},y^{i})\}_{i=1}^{k},\quad y^{i}=J(\boldsymbol{\mathrm{\theta}}^{i})+\epsilon^{i}$ and $\epsilon^{i}\sim\mathcal{N}(0,\sigma_{n}^{2})$, we then have
@@
where $\boldsymbol{\mathrm{\Theta}}=\left\{\boldsymbol{\mathrm{\theta}}^{i}\right\}_{i=1}^{k}$ and $K$ denotes the covariance matrix ($n\times n$ kernel matrix with elements $[K]_{(i,j)}=k(\boldsymbol{\mathrm{\theta}}_{i},\boldsymbol{\mathrm{\theta}}_{j})$) computed over the training inputs. The posterior distribution of $J(\boldsymbol{\mathrm{\theta}}^{\star})$ at a new input $\boldsymbol{\mathrm{\theta}}^{\star}$ is then given by
@@
$\begin{bmatrix}\mathbf{y}\\ J^{\star}\end{bmatrix}\sim\mathcal{N}\left(\begin{bmatrix}\mu(\boldsymbol{\mathrm{\Theta}})\\ \mu(\boldsymbol{\mathrm{\theta}}^{\star})\end{bmatrix},\begin{bmatrix}K+\sigma_{n}^{2}I&k(\boldsymbol{\mathrm{\theta}}^{\star},\boldsymbol{\mathrm{\Theta}})\\ k(\boldsymbol{\mathrm{\Theta}},\boldsymbol{\mathrm{\theta}}^{\star})&k(\boldsymbol{\mathrm{\theta}}^{\star},\boldsymbol{\mathrm{\theta}}^{\star})\end{bmatrix}\right).$
@@
This allows us to derive the predictive mean and variance for $J(\boldsymbol{\mathrm{\theta}}^{\star})$ as
$\mu(\boldsymbol{\mathrm{\theta}}^{\star})=k(\boldsymbol{\mathrm{\theta}}^{\star},\boldsymbol{\mathrm{\Theta}})(K+\sigma_{n}^{2}I)^{-1}\mathbf{y},$
$\sigma^{2}(\boldsymbol{\mathrm{\theta}}^{\star})=k(\boldsymbol{\mathrm{\theta}}^{\star},\boldsymbol{\mathrm{\theta}}^{\star})-k(\boldsymbol{\mathrm{\theta}}^{\star},\boldsymbol{\mathrm{\Theta}})(K+\sigma_{n}^{2}I)^{-1}k(\boldsymbol{\mathrm{\Theta}},\boldsymbol{\mathrm{\theta}}^{\star}).$
The updated GP models are then used to induce an acquisition function $\alpha(\boldsymbol{\mathrm{\theta}})$, which helps find the optimum of the unknown objective function $J(\boldsymbol{\mathrm{\theta}})$, that is
$\boldsymbol{\mathrm{\theta}}^{\star}=\arg\min_{\boldsymbol{\mathrm{\theta}}\in\mathcal{P}}\alpha(\boldsymbol{\mathrm{\theta}}).$ (37)
In this paper, we use the Expected Improvement (EI) as an acquisition function:
$\alpha_{\text{EI}}(\boldsymbol{\mathrm{\theta}})=\mathbb{E}[\min(J(\boldsymbol{\mathrm{\theta}})-J^{\star},0)],$ (38)
@@
where $J^{\star}$ is the best observed function value. This acquisition function can be computed using the properties of the Gaussian distribution as
@@
$\alpha_{\text{EI}}(\boldsymbol{\mathrm{\theta}})=-\left(\left(J^{\star}-\mu(\boldsymbol{\mathrm{\theta}})\right)\Phi(Z)+\sigma(\boldsymbol{\mathrm{\theta}})\phi(Z)\right),$ (39)
where $Z=\frac{J^{\star}-\mu(\boldsymbol{\mathrm{\theta}})}{\sigma(\boldsymbol{\mathrm{\theta}})}$, and $\Phi(Z)$ and $\phi(Z)$ denote the standard normal Cumulative Density Function (CDF) and Probability Density Function (PDF) of the standard normal distribution, respectively. As discussed previously, the aim is to leverage BO to learn a parametric distributed MPC so that local MPCs can deliver the corresponding optimal policies $\boldsymbol{\mathrm{\pi}}^{\boldsymbol{\mathrm{\theta}}_{i}^{\star}}$. Therefore, the BO-based learning method must be able to capture the optimal parameters $\boldsymbol{\mathrm{\theta}}^{\star}$. However, popular acquisition functions in the open literature have a greedy and myopic structure that selects the next best point based solely on its immediate impact without considering future rewards. This then causes the process to disregard the long-term effect of the next sample, potentially preventing the learning process from obtaining $\boldsymbol{\mathrm{\theta}}^{\star}$. To address this issue, non-myopic BO methods have been formulated in the MDP framework where they can frame the exploration-exploitation problem as a balance of immediate and future rewards [24, 25, 26].
Proposition 1 (BO as Finite-Horizon MDP).
Bayesian optimization (BO) with the Expected Improvement (EI) acquisition function can be formulated as a finite-horizon Markov Decision Process (MDP) of horizon $H$, where the optimal query point $\boldsymbol{\mathrm{\theta}}^{\star}$ minimizes the non-myopic acquisition function
$\tilde{\alpha}_{\text{EI}}(\boldsymbol{\mathrm{\theta}})=\alpha_{\text{EI}}(\boldsymbol{\mathrm{\theta}})\qquad\qquad\text{if}\quad H=1,$ (40)
$\tilde{\alpha}_{\text{EI}}(\boldsymbol{\mathrm{\theta}})=V^{H,\boldsymbol{\mathrm{\tau}}}\left(\mathcal{D}^{k}\right)\qquad\text{if}\quad H>1,$
with $V^{H,\boldsymbol{\mathrm{\tau}}}\left(\mathcal{D}^{k}\right)$ denoting the associated value function under policy $\boldsymbol{\mathrm{\tau}}:\mathcal{D}^{k}\rightarrow\boldsymbol{\mathrm{\theta}}^{k+1}$.
Proof.
Let us view the BO as an $H$-step Dynamic Programming (DP) by casting it as a finite-horizon MDP for which the state and action are $\mathcal{D}^{k}$ and $\boldsymbol{\mathrm{\theta}}^{k+1}$, respectively. Considering $y^{i}=J(\boldsymbol{\mathrm{\theta}}^{i})+\epsilon^{i}$ and $\epsilon^{i}\sim\mathcal{N}(0,\sigma_{n}^{2})$, the collected data is then a Markov process, and hence the optimal policy $\boldsymbol{\mathrm{\tau}}^{\star}:\mathcal{D}^{k}\rightarrow\boldsymbol{\mathrm{\theta}}^{\star,k+1}$ must satisfy the Bellman principle of optimality [27]. Let us define a stage cost function as
$L\left(\mathcal{D}^{\ell},\boldsymbol{\mathrm{\theta}}^{\ell+1}\right)=\min(y(\boldsymbol{\mathrm{\theta}}^{\ell+1})-y^{\star,\ell},0),$ (41)
where $y^{\star,\ell}$ denotes the minimum observed value in the set $\mathcal{D}^{\ell}$, i.e., $y^{\star,\ell}=\min\left\{y^{0},\ldots,y^{\ell}\right\}$. We then define the corresponding finite-horizon value function as
$V^{H,\boldsymbol{\mathrm{\tau}}}\left(\mathcal{D}^{k}\right)=$ $\mathbb{E}_{y}\left[\sum_{\ell=k}^{k+H}L\left(\mathcal{D}^{\ell},\boldsymbol{\mathrm{\tau}}^{\ell}\right)\right].$ (42)
The Bellman equation then reads as
$V^{H,\boldsymbol{\mathrm{\tau}}}\left(\mathcal{D}^{k}\right)=\mathbb{E}_{y}\left[L\left(\mathcal{D}^{k},\boldsymbol{\mathrm{\tau}}^{k}\right)\right]+\mathbb{E}_{y}\left[V^{H,\boldsymbol{\mathrm{\tau}}}\left(\mathcal{D}^{+}\right)\right].$ (43)
We observe that
$\mathbb{E}_{y}\left[L\left(\mathcal{D}^{k},\boldsymbol{\mathrm{\tau}}^{k}\right)\right]=\mathbb{E}_{y}\left[\min(y(\boldsymbol{\mathrm{\theta}}^{\ell+1})-y^{\star,\ell},0)\right]=\alpha_{\text{EI}}(\boldsymbol{\mathrm{\theta}}),$ (44)
such that the expected improvement acquisition function then satisfies the Bellman equation as
$V^{H,\boldsymbol{\mathrm{\tau}}}\left(\mathcal{D}^{k}\right)=\alpha_{\text{EI}}(\boldsymbol{\mathrm{\theta}})+\mathbb{E}_{y}\left[V^{H,\boldsymbol{\mathrm{\tau}}}\left(\mathcal{D}^{+}\right)\right].$ (45)
For $H>1$, we then obtain the optimal solution by minimizing the non-myopic acquisition function as
$\boldsymbol{\mathrm{\theta}}^{\star}=\boldsymbol{\mathrm{\tau}}\left(\mathcal{D}^{k}\right)=\operatorname*{arg\,min}_{\boldsymbol{\mathrm{\theta}}}V^{H,\boldsymbol{\mathrm{\tau}}}\left(\mathcal{D}^{k}\right).$ (46)
It is noted that the second term in (45) can be practically computed using Monte Carlo (MC) simulations as
$\mathbb{E}_{y}\left[V^{H,\boldsymbol{\mathrm{\tau}}}\left(\mathcal{D}^{+}\right)\right]\approx\frac{1}{M}\sum_{i=1}^{M}\sum_{\ell=k}^{k+H-1}\left(y^{\star,\ell}-y^{\ell+1}\right)^{+}.$ (47)